OpenStudy (anonymous):

What is the volume of a pyramid with base area B = 6 ft2 and height h = 13 ft? A. 26 ft3 B. 39 ft3 C. 52 ft3 D. 78 ft3

OpenStudy (anonymous):

General formula for the volume of a pyramid: $$ \text{Area of the base} \times \text{Height} \times \frac{1}{3} = \text{Volume of a Pyramid} $$

OpenStudy (anonymous):

its A 36 ft3

OpenStudy (anonymous):

The base of a square pyramid has a side length of 6 inches. The height of the pyramid is 12 inches. What is the volume of the pyramid? A. 72 in3 B. 36 in3 C. 144 in3 D. 108 in3

OpenStudy (anonymous):

ops sorry

OpenStudy (anonymous):

@Bailee and thanx

OpenStudy (anonymous):

The base of a square pyramid has a side length of 6 inches. The height of the pyramid is 12 inches. What is the volume of the pyramid? Okay so we know that the pyramid has a square base, and that all sides of a square are the same. To get the area of the base that is a square we do: $$ 6 \times 6 \times 6 \times 6 = \text{Area of base} $$ Since we already know the height, just sub in the values into the formula I gave above. :-)